OpenStudy (anonymous):

x+1(lessthan or = to >) 2x+5 what is the solution set?

OpenStudy (across):

\[x+1\leq2x+5?\]

OpenStudy (anonymous):

> this way

OpenStudy (turingtest):

\[x+1\ \ge2x+5\]?

OpenStudy (across):

\[x+1\geq2x+5!?\]

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

\[-4\ge x\]

OpenStudy (turingtest):

subtract like terms on each side

OpenStudy (anonymous):

thats not it though

OpenStudy (anonymous):

(-4,infiniti????)

OpenStudy (across):

\[x+1\geq2x+5\]\[1\geq x+5\]\[-4\geq x\]Therefore, the solution set is\[\{x\in\mathbb{R}:x\leq-4\}\]or similarly\[(-\infty,-4]\]

OpenStudy (turingtest):

so they are the same, it's just a notation thing.

OpenStudy (across):

Correct.